I am no good at word problems, is there anyone who can help me with the problem below?
The cumulative number of car accidents from 2000 to 2010 can be modeled by the quadratic expression C= 25x^2 + 550x +7,300, where x=1 corresponds to 2000, x= 2 corresponds to 2001, and so on until x = 11 corresponds to 2010.
Find the number of car accidents in 2004 ?
Thank you so much.

Found 2 solutions by jorel1380, stanbon:
Answer by jorel1380(3719)

x=1=2000
x=2=2001
x=3=2002
x=4=2003
x=5=2005
---------------
C=25(5)2+550(5)+7300
"=25x25+2750+7300
"=625+10050
C=10675= the number of car accidents in 2004.

Answer by stanbon(75887)

The cumulative number of car accidents from 2000 to 2010 can be modeled by the quadratic expression C= 25x^2 + 550x +7,300, where x=1 corresponds to 2000, x= 2 corresponds to 2001, and so on until x = 11 corresponds to 2010.
Find the number of car accidents in 2004 ?
---------------------------------
C= 25x^2 + 550x +7,300
---
C(5) = 25(5^2) + 550(5) + 7300
---
C(5) = 10,675
======================
Cheers,
Stan H.